OpenStudy (anonymous):

simplify (x^2 -36)/(x^2 + 3x -18) . please someone answer

OpenStudy (anonymous):

x2-36=(x+6)(x-6) x2+3x-18=x2+6x-3x+18=x(x+6)-3(x+6)=(x-3)(x+6) cancel (x+6) you get (x-6)/(x-3)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

hope you understood the process??